# Preference Elicitation for Multi-objective Combinatorial Optimization with Active Learning and Maximum Likelihood Estimation

### v1

@@ Abstract
Real-life combinatorial optimization problems often involve several conflicting objectives, such as price, product quality and sustainability. A computationally-efficient way to tackle multiple objectives is to aggregate them into a single-objective function, such as a linear combination. However, defining the weights of the linear combination upfront is hard; alternatively, the use of interactive learning methods that ask users to compare candidate solutions is highly promising. The key challenges are to generate candidates quickly, to learn an objective function that leads to high-quality solutions and to do so with few user interactions. We build upon the Constructive Preference Elicitation framework and show how each of the three properties can be improved: to increase the interaction speed we investigate using pools of (relaxed) solutions, to improve the learning we adopt Maximum Likelihood Estimation of a Bradley-Terry preference model; and to reduce the number of user interactions, we select the pair of candidates to compare with an ensemble-based acquisition function inspired from Active Learning. Our careful experimentation demonstrates each of these improvements: on a PC configuration task and a realistic multi-instance routing problem, our method selects queries faster, needs fewer queries and synthesizes higher-quality combinatorial solutions than previous CPE methods.
@@ 2 Related work
It aims to estimate the parameters of a utility function representing preferences. It relies on an incremental interactive approach of selecting a query and updating the parameter estimate based on the feedback received Pigozzi et al. (2016). Most approaches aim to select the best choice out of an explicit finite set of alternatives. As such, they are not constructive as we consider here. Examples include multi-attribute utility theory Braziunas and Boutilier (2007) and multi-criteria decision-making Martyn and Kadziński (2023); Guo et al. (2021). Moreover, these approaches are either Bayesian and thus computationally expensive to train Guo and Sanner (2010), or exact regret-based and not robust to inconsistent answers Boutilier (2013); Toffano et al. (2022). All these works are non-contextual as they consider single-instance problems.
@@ 3 Problem Statement
We consider a CO problem defined over a set of variables $\{y_{1},\ldots,y_{k}\}$, where each variable $y_{i}$ takes values from its corresponding domain $D_{i}$. The Cartesian product of the domains, $\mathcal{Y}=\prod_{i=1}^{k}D_{i}$, represents the space of all possible assignments. This space is restricted by a set of constraints, defining a feasible region $\mathcal{F}\subseteq\mathcal{Y}$. The quality of an assignment $y\in\mathcal{Y}$ (whether feasible or not) is evaluated using $n$ sub-objectives, represented by a function $\phi:\mathcal{Y}\rightarrow\mathds{R}^{n}$. As common in utility theory Keeney (1993), we assume sub-objectives can be aggregated into a utility function $u$ that captures the DM’s preferences over assignments. An assignment $y_{+}$ is preferred to $y_{-}$, denoted by $y_{+}\succ y_{-}$, iff $u(y_{+})>u(y_{-})$. We further assume the utility to be linear in the sub-objectives $\phi$:
$u(y)=\langle w,\phi(y)\rangle=\sum_{i}^{n}w_{i}\phi_{i}(y)$
@@
Formally, the set of preference observations to learn from will be $\{(p,(y_{1},y_{2}),a))\}$ with $y_{1}$ and $y_{2}$ two assignments of the instance $p$. The preference label is $a=1$ if $y_{1}\succ y_{2}$, $a=-1$ if $y_{2}\succ y_{1}$ and $a=0$ if the DM is indifferent. The data is collected from the DM one data point at a time, during the preference elicitation process. We aim to improve the overall interaction by selecting queries in real-time and minimizing the number of queries required for effective learning.
@@
Input: number of steps $T$, number of sub-objectives $n$, number of models $k$, number of clusters
@@
1:for $i\in\{1,\ldots,k\}$ do $\triangleright$ Initial weight estimate
@@
15:     for $i\in\{1,\ldots,k\}$ do $\triangleright$ Weight update
@@ 4.1 Active Learning-based Query Selection
In CPE, both uncertainty and solution quality matter. However, variance is a non-linear function, making it difficult for a solver to optimize. Identifying the solution with the highest variance in the pool is much simpler. Thus, the pre-computed pool enables us to propose an ensemble-based acquisition function that balances both uncertainty and solution quality. We propose to train an ensemble $W^{t}=\{W_{1}^{t},\ldots,W_{n}^{t}\}$ composed of $k$ weight vectors $W_{i}^{t}\in\mathds{R}^{n}$, all initialized differently following a normal distribution of mean 1 and variance 1 (line 2 in Algorithm 2). Uncertainty is measured by the variance $\sigma$ of the ensemble $W_{t}$ (at time step $t$) and solution quality by its mean $\mu$. It corresponds to the UCB Archetti and Candelieri (2019):
$\textrm{UCB}(W^{t},y)=(1-\gamma)\mu(y)+\gamma\sigma(y)$
@@
The two highest-scoring UCB solutions may be highly similar or even identical in sub-objective values. A DM is likely to be indifferent to such a query, which would not provide a training signal. To ensure more diversity, previous work in preference-based Reinforcement Learning Marta et al. (2023) proposed to cluster the samples in a (learned) latent representation space. In the case of CO, the sub-objectives $\phi:Y\rightarrow\mathds{R}^{n}$ can be interpreted as a latent representation of an assignment $y$. Therefore, we cluster the solution pool based on the sub-objective values, using k-means. At each iteration, two clusters are randomly chosen and the solution with the highest UCB is selected in each cluster (line 11 in Algorithm 2). The resulting pairwise query is both informative (thanks to UCB) and diverse (due to clustering). Our query selection is illustrated in Figure 1.
@@
The time complexity of a single query selection is the complexity of sorting the points of two clusters by UCB value, $O(2*m\log(m))$ with $m$ the size of the largest cluster. It does not depend on the complexity of the CO problem considered, resulting in high-efficiency even for NP-hard and large-scale problems. The computation time of the solutions is amortized in the pre-computation of a large pool of feasible or relaxed solutions, which can easily be parallelized.
@@ 4.2.2 A unified view
There is $\alpha^{t}=\mathds{1}_{u^{t}(y_{+})<u^{t}(y_{-})}$ for SP, $\alpha^{t}=1$ for PP/NCE and $\alpha^{t}=\mathcal{S}(u^{t}(y_{-})-u^{t}(y_{+}))$ for MLE. The update factor $\alpha^{t}$ ranges between 00 and $1$ and controls the amplitude of the gradient step amplitude. As plotted in Figure 2, MLE can be interpreted as a smoothed version of SP. When the difference in predicted utility between $y_{+}$ and $y_{-}$ is large, i.e., a clear preference is predicted, the update factor tends to 00 if the prediction is correct and is maximum ($1$) otherwise. When no clear preference is predicted , the MLE update factor is a function of the utility difference.
@@ 5.1 Experimental setting
We first consider the single-instance task of PC Configuration used by both baselines as a constructive variant of PC recommendation Guo and Sanner (2010). Each configuration is described by $7$ components – such as storage and CPU model – and the price. Sub-objectives are Boolean indicating if a component is selected plus the normalized price. We also consider the more realistic multi-instance task of Prize-Collecting Traveling Salesperson Problem (PC-TSP), previously used for structured prediction Véjar et al. (2024). Each node represents a city, which offers a prize if visited and costs a penalty $\rho$ if not. Each edge $(i,j)$ has $k=4$ values $v_{k}^{ij}$ that can be interpreted as distances, duration, fuel consumption and driver familiarity. The total penalty is a $5$th sub-objective. Mathematically, the CO problem objective is
@@
We replicate the same experimental protocol as baselines. Each experiment is repeated $20$ times, each with a different DM whose ground-truth utility is obtained by randomly sampling each weight. For the configuration task, DM weights were sampled from a normal distribution $\mathcal{N}(25,25/3)$ Teso et al. (2016) then $80\%$ were randomly set to 00 to create sparse preferences. For the PC-TSP, DM weights were sampled from a Dirichlet’s distribution of concentration parameter $100$ Véjar et al. (2024).
@@
The code will be made available publicly on GitHub. The experiments are implemented using CPMpy Guns (2019) and GurobiPy Gurobi Optimization, LLC (2024). MLE was trained both online and with batch training, the latter for $4$ epochs and a batch size of $4$. Parameters were tuned based on a separate set of $10$ DMs. After tuning, the learning rate for the configuration task is $2$; for PC-TSP it is $0.5$ for SP-online and MLE-online, $0.1$ for PP-online and $1$ for MLE-batch. We applied $100$ training steps like Teso et al. (2016); Dragone et al. (2018b). For the ensemble-based acquisition functions, an ensemble of $25$ models is used. Since the sub-objectives are Boolean for the Configuration problem, all configurations with the same number of identical components are equidistant, so no clustering is applied. For PC-TSP, the number of clusters used was $5$. For each problem, a pool of $10,000$ different random relaxed solutions was generated, by randomly selecting attributes (configuration task) or by randomly sampling node permutations (PC-TSP). For PC-TSP, varying-size circuits were generated (the circuit is completed when the warehouse node is reached).
@@ 5.3 Impact of query selection and weight update
#Q ()($\downarrow$)( ↓ ) is the number of queries to reach 10% regret, % DM ()($\uparrow$)( ↑ ) is the percentage of satisfied DMs. For regret ()($\downarrow$)( ↓ ), standard errors are given.
@@ 5.4 Ablations
We also verify the effect of using clustering to increase diversity (only applicable to PC-TSP). We here describe the results of that experiment: when no clustering was used, $98.3$% of selected pairs were similar, i.e., they would have belonged to the same cluster. This lack of diversity led to more indifferent answers: $8.0$%, vs $0.7$% with clustering. The quality of synthesized solutions also dropped, with a regret of $13.5$% (vs $0.5$%) and only $40\%$ (vs 100%) of DMs were satisfied.

### v2

@@ Abstract
Real-life combinatorial optimization problems often involve several conflicting objectives, such as price, product quality and sustainability. A computationally-efficient way to tackle multiple objectives is to aggregate them into a single-objective function, such as a linear combination. However, defining the weights of the linear combination upfront is hard; alternatively, the use of interactive learning methods that ask users to compare candidate solutions is highly promising. The key challenges are to generate candidates quickly, to learn an objective function that leads to high-quality solutions and to do so with few user interactions. We build upon the Constructive Preference Elicitation (CPE) framework and show how each of the three properties can be improved: to increase the interaction speed we investigate using pools of (relaxed) solutions, to improve the learning we adopt Maximum Likelihood Estimation of a Bradley-Terry preference model; and to reduce the number of user interactions, we select the pair of candidates to compare with an ensemble-based acquisition function inspired from Active Learning. Our careful experimentation demonstrates each of these improvements: on a PC configuration task and a realistic multi-instance routing problem, our method selects queries faster, needs fewer queries and synthesizes higher-quality combinatorial solutions than previous CPE methods.
@@ 2 Related Work
It aims to estimate the parameters of a utility function representing preferences. It relies on an incremental interactive approach of selecting a query and updating the parameter estimate based on the feedback received Pigozzi et al. (2016). Most approaches aim to select the best choice out of an explicit finite set of alternatives.As such, they are not constructive as we consider here. Examples include multi-attribute utility theory Braziunas and Boutilier (2007) and multi-criteria decision-making Martyn and Kadziński (2023); Guo et al. (2021). Moreover, these approaches are either Bayesian and thus computationally expensive to train Guo and Sanner (2010), or exact regret-based and not robust to inconsistent answers Boutilier (2013); Toffano et al. (2022). All these works are non-contextual as they consider single-instance problems. Some non-constructive approaches use more sophisticated aggregation function than a linear utility, such as Choquet integral Herin et al. (2024); Vernerey et al. (2024).
@@ 3 Problem Statement
We consider a CO problem defined over a set of variables $\{y_{1},\ldots,y_{l}\}$, where each variable $y_{i}$ takes values from its corresponding domain $D_{i}$. The Cartesian product of the domains, $\mathcal{Y}=\prod_{i=1}^{l}D_{i}$, represents the space of all possible assignments. This space is restricted by a set of constraints, defining a feasible region $\mathcal{F}\subseteq\mathcal{Y}$. The quality of an assignment $y\in\mathcal{Y}$ (whether feasible or not) is evaluated using $n$ sub-objectives, represented by a function $\phi:\mathcal{Y}\rightarrow\mathds{R}^{n}$. As common in utility theory Keeney (1993), we assume sub-objectives can be aggregated into a utility function $u$ that captures the DM’s preferences over assignments. An assignment $y_{+}$ is preferred to $y_{-}$, denoted by $y_{+}\succ y_{-}$, iff $u(y_{+})>u(y_{-})$. We further assume the utility to be linear in the sub-objectives $\phi$:
$u(y)=\langle w,\phi(y)\rangle=\sum_{i=1}^{n}w_{i}\phi_{i}(y)$
@@
Formally, the set of preference observations to learn from will be $\{(p,(y_{1},y_{2}),a))\}$ with $y_{1}$ and $y_{2}$ two assignments of the instance $p$. The preference label is $a=1$ if $y_{1}\succ y_{2}$, a=1a=-1 if $y_{2}\succ y_{1}$ and $a=0$ if the DM is indifferent. The data is collected from the DM one data point at a time, during the preference elicitation process. We aim to improve the overall interaction by selecting queries in real-time and minimizing the number of queries required for effective learning.
@@
Input: number of steps $T$, number of sub-objectives $n$, number of models $m$, number of clusters
@@
1:for $i\in\{1,\ldots,m\}$ do $\triangleright$ Initial weight estimate
@@
15:  for $i\in\{1,\ldots,m\}$ do $\triangleright$ Weight update
@@ 4.1 Active Learning-based Query Selection
In CPE, both uncertainty and solution quality matter. However, variance is a non-linear function, making it difficult for a solver to optimize. Identifying the solution with the highest variance in the pool is much simpler. Thus, the pre-computed pool enables us to propose an ensemble-based acquisition function that balances both uncertainty and solution quality. We propose to train an ensemble $W^{t}=\{W_{1}^{t},\ldots,W_{m}^{t}\}$ composed of $m$ weight vectors $W_{i}^{t}\in\mathds{R}^{n}$, all initialized differently following a normal distribution of mean 1 and variance 1 (line 2 in Algorithm 2). Uncertainty is measured by the variance $\sigma_{W^{t}}$ of the ensemble $W^{t}$ (at time step $t$) and solution quality by its mean $\mu_{W^{t}}$. It corresponds to the UCB Archetti and Candelieri (2019):
$\textrm{UCB}(W^{t},y)=(1-\gamma)\mu_{W^{t}}(y)+\gamma\sigma_{W^{t}}(y)$
@@
The two highest-scoring UCB solutions may be highly similar or even identical in sub-objective values. A DM is likely to be indifferent to such a query, which would not provide a training signal. To ensure more diversity, previous work in preference-based Reinforcement Learning Marta et al. (2023) proposed to cluster the samples in a (learned) latent representation space. In the case of CO, the sub-objectives $\phi:Y\rightarrow\mathds{R}^{n}$ can be interpreted as a latent representation of an assignment $y$. Therefore, we cluster the solution pool based on the sub-objective values, using k-means++.At each iteration, two clusters are randomly chosen and the solution with the highest UCB is selected in each cluster (line 11 in Algorithm 2). The resulting pairwise query is both informative (thanks to UCB) and diverse (due to clustering). Our query selection is illustrated in Figure 1.
@@
The time complexity of a single query selection is the complexity of sorting the points of two clusters by UCB value, $O(2*q\log(q))$ with $q$ the size of the largest cluster. It does not depend on the complexity of the CO problem considered, resulting in high-efficiency even for NP-hard and large-scale problems. The computation time of the solutions is amortized in the pre-computation of a large pool of feasible or relaxed solutions, which can easily be parallelized.
@@ 4.2.2 A unified view
There is $\alpha^{t}=\mathds{1}_{u^{t}(y_{+})<u^{t}(y_{-})}$ for SP, $\alpha^{t}=1$ for PP/NCE and $\alpha^{t}=\mathcal{S}(u^{t}(y_{-})-u^{t}(y_{+}))$ for MLE. The update factor $\alpha^{t}$ ranges between 0 and $1$ and controls the amplitude of the gradient step amplitude. As plotted in Figure 2, MLE can be interpreted as a smoothed version of SP. When the difference in predicted utility between $y_{+}$ and $y_{-}$ is large, i.e., a clear preference is predicted, the update factor tends to 0 if the prediction is correct and is maximum ($1$) otherwise. When no clear preference is predicted , the MLE update factor is a function of the utility difference.
@@ 5.1 Experimental Setting
We first consider the single-instance task of PC Configuration used by both baselines as a constructive variant of PC recommendation Guo and Sanner (2010). Each configuration is described by $7$ components – such as storage and CPU model – and the price. Sub-objectives are Boolean ($77$ in total) indicating if a component is selected, plus the normalized price. We also consider the more realistic multi-instance task of Prize-Collecting Traveling Salesperson Problem (PC-TSP), previously used for structured prediction Véjar et al. (2024). Each node represents a city, which offers a prize if visited and costs a penalty $\rho$ if not. Each edge $(i,j)$ has $k=4$ values $v_{k}^{ij}$ that can be interpreted as distances, duration, fuel consumption and driver familiarity. The total penalty is a $5$th sub-objective. Mathematically, the CO problem objective is
@@
We replicate the same experimental protocol as baselines. Each experiment is repeated $20$ times, each with a different DM whose ground-truth utility is obtained by randomly sampling each weight. For the configuration task, DM weights were sampled from a normal distribution $\mathcal{N}(25,25/3)$ Teso et al. (2016) then $80\%$ were randomly set to 0 to create sparse preferences. For the PC-TSP, DM weights were sampled from a Dirichlet’s distribution of concentration parameter $100$ Véjar et al. (2024).
@@
The experiments are implemented using CPMpy Guns (2019) and GurobiPy Gurobi Optimization, LLC (2024). MLE was trained both online and with batch training, the latter for $4$ epochs and a batch size of $4$. Parameters were tuned based on a separate set of $10$ DMs. After tuning, the learning rate for the configuration task is $2$; for PC-TSP it is $0.5$ for SP-online and MLE-online, $0.1$ for PP-online and $1$ for MLE-batch. We applied $100$ training steps like Teso et al. (2016); Dragone et al. (2018b). For the ensemble-based acquisition functions, an ensemble of $m=25$ models is used. Since the sub-objectives are Boolean for the Configuration problem, all configurations with the same number of identical components are equidistant, so no clustering is applied. For PC-TSP, the number of clusters used was $5$. For each problem, a pool of $10,000$ different random relaxed solutions was generated, by randomly selecting attributes (configuration task) or by randomly sampling node permutations (PC-TSP). For PC-TSP, varying-size circuits were generated (the circuit is completed when the warehouse node is reached). The code is available at: https://github.com/mdefresne/CPE_pool.
@@ 5.3 Impact of Query Selection and Weight Update
#Q $(\downarrow)$ is the number of queries to reach 10% regret, % DM $(\uparrow)$ is the percentage of satisfied DMs. For regret $(\downarrow)$, standard errors are given.
@@ 5.4 Ablations
We also verify the effect of using clustering to increase diversity (only applicable to PC-TSP). We here describe the results of that experiment: when no clustering was used, 98.398.3% of selected pairs were similar, i.e., they would have belonged to the same cluster. This lack of diversity led to more indifferent answers: 8.08.0%, vs 0.70.7% with clustering. The quality of synthesized solutions also dropped, with a regret of 13.513.5% (vs 0.50.5%) and only $40\%$ (vs 100%) of DMs were satisfied.
@@ 6 Conclusion
Acknowledgments
This research received funding from the European Research Council (ERC) under the EU Horizon 2020 research and innovation programme (Grant No 101002802, CHAT-Opt). Jayanta Mandi is supported by the Research Foundation-Flanders (FWO) project G0G3220N.